# Relativistic stars in $f(Q)$-gravity: Exact analytic solution for the power-law case $f(Q)=Q+b\>Q^{\nu}$

### v2

@@ IV The model under consideration
Figure 1: The radial (continuous lines) and tangential (dashed lines) sound speeds as functions of the star radius. The theory considered has $\nu=2$. The red lines correspond to $\alpha=0.1$ and the blue ones to $\alpha=0.01$. For the rest of the parameters we have $r_{1}=r_{2}=10$ km and $Q_{0}=0.01$ km-2.
Figure 2: The mass function in the interior of the star. The power index characterizing the theory is $\nu=2$. The red lines correspond to $\alpha=0.1$ and the blue ones to $\alpha=0.01$. For both curves $r_{1}=r_{2}=10$ km and $Q_{0}=0.01$ km-2.
Figure 3: Normalized energy density (dotted curves) and pressures in both directions (radial: solid curves, tangential: dashed curves) versus radial coordinate assuming $\alpha=0.01$ in blue color and $\alpha=0.1$ in red color. All three quantities are positive throughout the star, while at the same time $\rho>p_{r,t}$.
Figure 4: Relativistic adiabatic index, $\Gamma$, in both directions (radial: solid, tangential: dashed) versus radial coordinate considering $\alpha=0.01$ in blue color and $\alpha=0.1$ in red color. The horizontal line in black represents the Newtonian value $4/3$.
For the sake of comparison, we report here the stellar mass and radius for the cases of GR ($\alpha=0$) as well as the two numerical values for $\alpha=0.01,0.1$ for the $f(Q)$ gravity considered here. They are as follows:
$R=11.21km,\;\;\;\;\;M=2.70M_{\odot},\;\;\;\;\;\alpha=0$ (56)
$R=11.24km,\;\;\;\;\;M=2.71M_{\odot},\;\;\;\;\;\alpha=0.01$ (57)
$R=11.53km,\;\;\;\;\;M=2.81M_{\odot},\;\;\;\;\;\alpha=0.1$ (58)
@@ V Conclusions
We exploited the existence of this known vacuum solution to use it as the exterior metric in our problem. For the interior of the star, we encountered first the connection which is compatible with the field equations induced by a space-time given by line element (17). It is important to note that this connection can be used in this context with any $f(Q)$ theory. In our investigation we used a power-law $f(Q)$ function (38) and, under the ansatz (39), we determined the physical properties of the respective fluid. We determined variables of the parameters, which give a reasonable behavior for the energy density and the two pressures. We identified the radius $r=\mathcal{R}$, where we assume that $Q(\mathcal{R})=\text{const}$. for $r\geq\mathcal{R}$. In this setting we calculated the explicit expression for the mass function.
Additionally, we analyzed the radial and tangential sound speeds near the singularity. For values $\nu>2$ of the power index characterizing the $f\left(Q\right)$-theory, the variation of the sound speeds is nearly independent of $\nu$. These findings contribute to a deeper understanding of the role of nonmetricity in gravitational models and support the viability of $f(Q)$ gravity in describing compact objects. In a future work we plan to study the definition for the Tolman - Oppenheimer - Volkov formula within the symmetric teleparallel theory of gravity.
@@ Acknowledgements.
AG was supported by Proyecto Fondecyt Regular 1240247. AP thanks the support of Vicerrectoría de Investigación y Desarrollo Tecnológico (Vridt) at Universidad Católica del Norte through Núcleo de Investigación Geometría Diferencial y Aplicaciones, Resolución Vridt No - 096/2022 and Resolución Vridt No - 098/2022. AP was partially supported by Proyecto Fondecyt Regular 2024, Folio 1240514, Etapa 2024. AP thanks the Universidad de La Frontera for the hospitality provided when this work was carried out.

### v3

@@ IV.2 Physical relevance of the solution
Figure 1: The radial (solid lines) and tangential (dashed lines) sound speeds as functions of the stellar radius considering $\nu=2$. The red lines correspond to $\alpha=0.10$ and the blue ones to $\alpha=0.01$. For the rest of the parameters we have assumed $r_{1}=r_{2}=10.00$ km and $Q_{0}=0.01$ km-2.
Figure 2: The mass function in the interior of the star. The power index characterizing the theory is $\nu=2$. The red lines correspond to $\alpha=0.10$ and the blue ones to $\alpha=0.01$. For both curves $r_{1}=r_{2}=10.00$ km and $Q_{0}=0.01$ km-2.
Figure 3: Normalized energy density (dotted curves) and pressures in both directions (radial: solid curves, tangential: dashed curves) versus radial coordinate assuming $\alpha=0.01$ in blue color and $\alpha=0.10$ in red color. All three quantities are positive throughout the star, while at the same time $\rho>p_{r,t}$.
Figure 4: Relativistic adiabatic index, $\Gamma$, in both directions (radial: solid, tangential: dashed) versus radial coordinate considering $\alpha=0.01$ in blue color and $\alpha=0.10$ in red color. The horizontal line in black represents the Newtonian value $4/3$.
For the sake of comparison, we report here the stellar mass and radius as well as the factor of compactness, $C=M/R$, for the cases of GR ($\alpha=0.00$) as well as the two numerical values for $\alpha=0.01,\>0.10$ for the $f(Q)$ gravity considered here. They are as follows:
$R=11.21km,\;\;\;\;M=2.70M_{\odot},\;\;\;\;C=M/R=0.358,\;\;\;\;\alpha=0.00$ (58)
$R=11.24km,\;\;\;\;M=2.71M_{\odot},\;\;\;\;C=M/R=0.359,\;\;\;\;\alpha=0.01$ (59)
$R=11.53km,\;\;\;\;\;M=2.81M_{\odot},\;\;\;\;C=M/R=0.363,\;\;\;\;\;\alpha=0.10$ (60)
Clearly, an increasing coupling $\alpha$ implies somewhat larger and more massive stars, while at the same time the objects become more and more compact.
@@
The results show that as the coupling $\alpha$ increases, both the energy density and the pressures decrease faster with $r$. Moreover, the impact of $\alpha$ on the relativistic adiabatic index is not considerable. Finally, as far as the sound speeds are concerned, the increase of the parameter $\alpha$ shifts the curves downwards at low $r$, although as we approach the stellar radius the speed of sound becomes larger than the one corresponding to a lower coupling $\alpha$.
Before we conclude our work, a comment is in order here. As we have already mentioned, the primary objective of this study was to establish the mathematical and conceptual consistency of stellar interior solutions in the context of $f(Q)$ gravity, rather than to propose a definitive astrophysical model. What is more, since the exact analytic solution obtained here was based on a metric ansatz without adopting a concrete equation-of-state, it should be understood as an effective model rather than a microphysically realistic description. For those reasons it does not make sense to generate the M-R relationships and compare against current astrophysical constraints. This kind of analysis is performed in order to constrain realistic and well motivated equations-of-state derived using the nuclear physics, particle physics and statistical physics we know. The solution presented here is seen to describe realistic astrophysical configurations such as: neutron stars, quark stars or hyperon stars. The only limitation is the lack of a concrete equation of state, and so a priori we do not know which class of these stars this solution represents.
The deviation of our solution from the one of GR is quantitative as the numerical values of the stellar masses and radii suggest, and this could have considerable impact on other observables such as the frequencies of oscillation of pulsating stars, tidal Love numbers, etc.
@@ V Conclusions
We exploited the existence of this known vacuum solution to use it as the exterior metric in our problem. For the interior of the star, we encountered first the connection which is compatible with the field equations induced by a spacetime given by line element (17). It is important to note that this connection can be used in this context with any $f(Q)$ theory. In our investigation we used a power-law $f(Q)$ function (40) and, under the ansatz (41), we determined the physical properties of the respective fluid. We determined variables of the parameters, which give a reasonable behavior for the energy density and the two pressures. We identified the radius $r=\mathcal{R}$, where we assume that $Q(\mathcal{R})=\text{const}$. for $r\geq\mathcal{R}$. In this setting we calculated the explicit expression for the mass function. In this context we derived an exact solution to demonstrate the rigor of the procedure. A more physical approach would involve specifying an equation-of-state rather than imposing an ansatz for the metric. However, the resulting solution would then necessarily be numerical. Furthermore, one may go beyond spherical symmetry and add a non-vanishing angular momentum to study properties of rotating stars. We plan to extend this work by investigating astrophysical aspects through the consideration of different equations-of-state in various $f(Q)$ scenarios. Moreover, it would be interesting to study photon orbits and see how they deviate from the null geodesics within GR. We hope to be able to address some of those important topics in the future.
In the model under consideration, we analyzed the radial and tangential sound speeds near the singularity. For values $\nu>2$ of the power index characterizing the $f\left(Q\right)$-theory, the variation of the sound speeds is nearly independent of $\nu$. These findings contribute to a deeper understanding of the role of nonmetricity in gravitational models and support the viability of $f(Q)$ gravity in describing compact objects. In a future work we plan to study the definition for the Tolman - Oppenheimer - Volkov formula within the symmetric teleparallel theory of gravity.
@@ Acknowledgements.
We thank the anonymous reviewers for useful comments and suggestions. A. G was supported by Proyecto Fondecyt Regular 1240247. A. P thanks the support of Vicerrectoría de Investigación y Desarrollo Tecnológico (Vridt) at Universidad Católica del Norte through Núcleo de Investigación Geometría Diferencial y Aplicaciones, Resolución Vridt No - 096/2022 and Resolución Vridt No - 098/2022. A. P was partially supported by Proyecto Fondecyt Regular 2024, Folio 1240514, Etapa 2024. A. P thanks the Universidad de La Frontera for the hospitality provided when this work was carried out.